# The most symmetric smooth cubic surface over a finite field of characteristic $2$

### v1

@@ 5. Cubic surfaces
Let $S$ be a smooth cubic surface over $\mathbf{F}$ such that the image $\Gamma$ of the absolute Galois group $\mathrm{Gal}(\mathbf{F}^{sep}/\mathbf{F})$ in $W(\mathrm{E}_{6})$ is isomorphic to $\mathbb{Z}/2\mathbb{Z}.$ Let $G$ be a subgroup of order $5$ in $\mathrm{Aut}(S).$ Then there is a quintuple of skew lines in quadratic position which are defined over $\mathbf{F}$ such that they form an orbit under the action of $G.$
@@
By Corollary 4.6 the image of the Galois group is generated by an element lying in the conjugacy class $A_{1}.$ So up to conjugation this element is equal to $\sigma$ defined in Remark 5.5. The generator $\gamma$ of $\Gamma$ and the generator $g$ of $G$ are commute. As by Lemma 4.3 there is a unique element of order $10$ up to conjugation in $W(\mathrm{E}_{6})$, we get that the element $\gamma g$ is unique up to conjugation in $W(\mathrm{E}_{6}).$ Therefore, without loss of generality we can assume that $r=\gamma g=\sigma\cdot(23456).$ Thus, $\gamma=r^{5}$ and $g=r^{2}.$
By the definition of $\sigma$ all the lines
$L_{12},\;L_{13},\;L_{14},\;L_{15},\;L_{16}$
are defined over $\mathbf{F}.$ By Lemma 5.8 they are in general position. As up to conjugation the group $G$ is generated by $g,$ we get the desired.
@@
Let $S$ be a smooth cubic surface over a finite field $\mathbf{F}$ such that the image $\Gamma$ of the absolute Galois group $\mathrm{Gal}(\overline{\mathbf{F}}/\mathbf{F})$ in $W(\mathrm{E}_{6})$ is isomorphic to $\mathbb{Z}/2\mathbb{Z}.$ Let $G$ be a subgroup of order $5$ in $\mathrm{Aut}(S).$ Then there is a $\mathbb{Z}/5\mathbb{Z}$-equivariant blowdown
$\pi:S\to Q,$
where $Q\simeq\mathrm{R}_{\mathbf{L}/\mathbf{F}}(\mathbb{P}^{1})$ such that $\mathbf{F}\subset\mathbf{L}$ is a separable quadratic extension of $\mathbf{F}.$
@@
This directly follows from Lemma 4.9 since $\Gamma$ commutes with the automorphism group of $S.$
@@
Assume the contrary. Then all exceptional lines on $S$ are defined over $\mathbb{F}_{2^{2k+1}}.$ By Lemma 5.9 there is $\mathbb{Z}/5\mathbb{Z}$-equivariant blowdown to $\mathbb{P}^{1}\times\mathbb{P}^{1}.$ However, by Lemma 2.3 the order of the group $\mathrm{Aut}(\mathbb{P}^{1}\times\mathbb{P}^{1})$ is not divisible by $5$
@@ 6. Proof of the main results
By Lemma 5.2 the automorphism group of a smooth cubic surface lies in $\mathrm{PSU}_{4}(\mathbb{F}_{2}).$
Let us consider smooth cubic surfaces over $\mathbb{F}_{4^{k}}.$ By Theorem 1.2 the automorphism group of the Fermat cubic surface over $\mathbb{F}_{4}$ is isomorphic to $\mathrm{PSU}_{4}(\mathbb{F}_{2}).$ This means that the automorphism group of maximal order of a smooth cubic surface over $\mathbb{F}_{4^{k}}$ is isomorphic to $\mathrm{PSU}_{4}(\mathbb{F}_{2}).$
@@
Let $S$ be a smooth cubic surface over $\mathbb{F}_{2^{2k+1}}$ such that its automorphism group is of maximal order among smooth cubic surfaces over $\mathbb{F}_{2^{2k+1}}.$ Then by Theorem 1.3 we have $|\mathrm{Aut}(S)|\geqslant 720.$ By Lemmas 4.7 and 5.2 we get $3$ candidates for being the largest automorphism group of a smooth cubic surface $S$:
@@
Let us prove assertion (i). Since $\Gamma$ is trivial, all lines on $S_{\overline{\mathbf{F}}}$ are defined over $\mathbf{F}.$ By Lemma 5.9 there are $5$ skew lines in quadratic position such that they form an orbit under the action of $\mathbb{Z}/5\mathbb{Z}\subset\mathrm{Aut}(S).$ By Corollary 5.10 we can blow them down $\mathbb{Z}/5\mathbb{Z}$-equivariantly and get $\mathbb{P}^{1}\times\mathbb{P}^{1}.$ The blowdown gives $5$ points in general position on $\mathbb{P}^{1}\times\mathbb{P}^{1}$ by Remark 5.7. By Propositions 3.5 and 3.6 we get that the set of these $5$ points is unique up to conjugation. Therefore, a smooth cubic surface such that its automorphism group contains $\mathbb{Z}/5\mathbb{Z}$ and $\Gamma$ is trivial is unique up to isomorphism. By Theorem 1.2 we get that the automorphism group of the Fermat cubic surface contains $\mathbb{Z}/5\mathbb{Z}.$ Thus, $S$ is isomorphic to the Fermat cubic surface.
Now let us prove assertion (ii). By Lemma 5.11 there are $5$ skew lines in quadratic position such that they form an orbit under the action of $\mathbb{Z}/5\mathbb{Z}\subset\mathrm{Aut}(S).$ By Corollary 5.12 we can blow them down $\mathbb{Z}/5\mathbb{Z}$-equivariantly and get the Weil restriction of scalars $Q\simeq\mathrm{R}_{\mathbf{L}/\mathbf{F}}(\mathbb{P}^{1}),$ where $\mathbf{F}\subset\mathbf{L}$ is a quadratic extension of $\mathbf{F}.$ The blowdown gives us $5$ points in general position on $Q$ by Remark 5.7. By Proposition 3.7 we get that the set of these $5$ points is unique up to automorphism. Therefore, a smooth cubic surface such that its automorphism group contains $\mathbb{Z}/5\mathbb{Z}$ and $\Gamma\simeq\mathbb{Z}/2\mathbb{Z}$ is unique up to isomorphism. By Theorem 1.3 we get that the automorphism group of the cubic surface (1.1) contains $\mathbb{Z}/5\mathbb{Z}.$ Thus, $S$ is isomorphic to the cubic surface (1.1).
@@ 7. Fermat cubic surface over fields of characteristic $2$
First of all, note that we have
(7.1) $\mathbb{Z}/2\mathbb{Z}\times\mathfrak{S}_{4}\subseteq\mathrm{Aut}(\mathcal{F}),$
where $\mathfrak{S}_{4}$ permutes the coordinates and $\mathbb{Z}/2\mathbb{Z}$ is generated by the element
(7.2) $h=\begin{pmatrix}0&1&1&1\\ 1&0&1&1\\ 1&1&0&1\\ 1&1&1&0\end{pmatrix}.$
From the equations of lines on the Fermat cubic surface we get that the Galois group of the field of definition of all lines on $\mathcal{F}$ is isomorphic to $\mathbb{Z}/2\mathbb{Z}.$ The generator $g$ of this group maps $\omega$ to $\omega^{2},$ where $\omega\in\overline{\mathbf{F}}$ is a non-trivial cube root of unity. So again from the equations of lines on the Fermat cubic surface one can see that the element $g$ fixes $3$ lines, and interchanges the lines in the six pairs of intersecting lines and in the six pairs of non-intersecting lines. By this description of the element from Lemma 5.3 we see that the element $g$ in the Weyl group $W(\mathrm{E}_{6})$ lies in the conjugacy class $A_{1}^{3}.$ By Lemma 4.3 the order of its centralizer is $96$ By the same lemma the determinant of $g$ is equal to $-1.$ As both $g$ and $\mathrm{Aut}(\mathcal{F})$ lie in the centralizer of $g$ and the element $g$ does not lie in $\mathrm{Aut}(\mathcal{F}),$ we have the isomorphism in (7.1) by the comparison of the orders.
@@
By Theorem 1.4(ii) the automorphism group of the cubic (1.1) is isomorphic to $\mathfrak{S}_{6},$ while by Proposition 1.6 the automorphism of the Fermat cubic surface is isomorphic to $\mathbb{Z}/2\mathbb{Z}\times\mathfrak{S}_{4}.$ This proves assertion (i).
The isomorphism between (1.1) and the Fermat cubic surface over $\mathbb{F}_{4}$ is given by the collineation
@@
where $\omega\in\mathbb{F}_{4}$ is a non-trivial cube root of unity, which maps the cubic surface (1.1) to the Fermat cubic surface. This proves assertion (ii).

### v2

@@ 5. Cubic surfaces
Let $S$ be a smooth cubic surface over $\mathbf{F}$ such that the image $\Gamma$ of the absolute Galois group $\mathrm{Gal}(\mathbf{F}^{sep}/\mathbf{F})$ in $W(\mathrm{E}_{6})$ is isomorphic to $\mathbb{Z}/2\mathbb{Z}.$ Let $G$ be a subgroup of order $5$ in $\mathrm{Aut}(S).$ Then there is a quintuple of skew lines in quadratic position such that they form an orbit under the action of $G.$
@@
Let $\gamma$ be a generator of $\Gamma$ and $g$ be a generator of $G.$ Since $\gamma$ and $g$ are commute in $W(\mathrm{E}_{6}),$ we can consider the group $D=G\times\Gamma\simeq\mathbb{Z}/10\mathbb{Z}.$ By Lemma 5.8(ii) there is an orbit of length $2$ of skew lines $l_{1}$ and $l_{2}$ on $S_{\overline{\mathbf{F}}}$ under the action of $D$ on the configuration of lines on $S_{\overline{\mathbf{F}}}.$ Therefore, this orbit under the action of $D$ is also the orbit of length $2$ under the action of $\Gamma.$ By Lemma 5.9 there are exactly $5$ lines on $S,$ say $\mathcal{L}_{1},\mathcal{L}_{2},\mathcal{L}_{3},\mathcal{L}_{4},\mathcal{L}_{5},$ which intersect both $l_{1}$ and $l_{2}.$ Since this quintuple of lines is unique we get that any $\mathcal{L}_{i}$ is $\Gamma$-invariant, which means that all of them are defined over $\mathbf{F}.$ Moreover, this quintuple is $G$-orbit of length $5,$ since otherwise $G$ acts trivially on $S.$
Let us blow down this quintuple. We get a del Pezzo surface $S^{\prime}$ of degree $8$ Assume that $S^{\prime}$ is a blowup of $\mathbb{P}^{2}$ in a point. Then there is a line on $S$ which is invariant under the action of $D.$ However, this contradicts Lemma 5.8(i). So we get that $S^{\prime}_{\overline{\mathbf{F}}}\simeq\mathbb{P}^{1}_{\overline{\mathbf{F}}}\times\mathbb{P}^{1}_{\overline{\mathbf{F}}}.$
@@
Let $S$ be a smooth cubic surface over a finite field $\mathbf{F}$ such that the image $\Gamma$ of the absolute Galois group $\mathrm{Gal}(\mathbf{F}^{sep}/\mathbf{F})$ in $W(\mathrm{E}_{6})$ is isomorphic to $\mathbb{Z}/2\mathbb{Z}.$ Let $G$ be a subgroup of order $5$ in $\mathrm{Aut}(S).$ Then there is a $\mathbb{Z}/5\mathbb{Z}$-equivariant blowdown $\pi:S\to Q,$ where $Q\simeq\mathrm{R}_{\mathbf{L}/\mathbf{F}}(\mathbb{P}^{1})$ such that $\mathbf{F}\subset\mathbf{L}$ is a separable quadratic extension of $\mathbf{F}.$
@@
This directly follows from Lemma 4.8, since $\Gamma$ commutes with the automorphism group of $S.$
@@
Assume the converse. Then all exceptional lines on $S_{\overline{\mathbb{F}}_{2}}$ are defined over the field $\mathbb{F}_{2^{2k+1}}.$ By Lemma 5.11 we get that there is $\mathbb{Z}/5\mathbb{Z}$-equivariant blowdown to $\mathbb{P}^{1}\times\mathbb{P}^{1}.$ However, by Lemma 2.3 the order of the group $\mathrm{Aut}(\mathbb{P}^{1}\times\mathbb{P}^{1})$ is not divisible by $5$
@@ 6. Proof of the main results
The inequality follows from Lemma 5.2. Let us consider smooth cubic surfaces over $\mathbb{F}_{4^{k}}.$ By Theorem 1.2 the automorphism group of the Fermat cubic surface over $\mathbb{F}_{4}$ is isomorphic to $\mathrm{PSU}_{4}(\mathbb{F}_{2}).$ This means that the automorphism group of maximal order of a smooth cubic surface over $\mathbb{F}_{4^{k}}$ is isomorphic to $\mathrm{PSU}_{4}(\mathbb{F}_{2}).$
@@
Let $S$ be a smooth cubic surface over $\mathbb{F}_{2^{2k+1}}$ such that its automorphism group is of maximal order among smooth cubic surfaces over $\mathbb{F}_{2^{2k+1}}.$ Then by Theorem 1.3 we have $|\mathrm{Aut}(S)|\geqslant 720.$ By [7, Theorem 1.1] this means that
$\mathrm{Aut}(S_{\overline{\mathbb{F}}_{2}})\simeq\mathrm{PSU}_{4}(\mathbb{F}_{2}).$
Therefore, the group $\mathrm{Aut}(S)$ is a subgroup in $\mathrm{PSU}_{4}(\mathbb{F}_{2}).$ By Lemmas 4.6 and 5.2 we get $3$ candidates for being the largest automorphism group of a smooth cubic surface $S$:
@@
Let us prove assertion (i). Since $\Gamma$ is trivial, all lines on $S_{\overline{\mathbf{F}}}$ are defined over $\mathbf{F}.$ By Lemma 5.11 there are $5$ skew lines in quadratic position such that they form an orbit under the action of $\mathbb{Z}/5\mathbb{Z}\subset\mathrm{Aut}(S).$ By Corollary 5.12 we can blow them down $\mathbb{Z}/5\mathbb{Z}$-equivariantly and get $\mathbb{P}^{1}\times\mathbb{P}^{1}.$ By Remark 5.5 the blowdown gives $5$ points in general position on $\mathbb{P}^{1}\times\mathbb{P}^{1}.$ By Propositions 3.5 and 3.6 we get that the set of these $5$ points is unique up to conjugation. Therefore, a smooth cubic surface such that its automorphism group contains $\mathbb{Z}/5\mathbb{Z}$ and $\Gamma$ is trivial is unique up to isomorphism. By Theorem 1.2 we get that the automorphism group of the Fermat cubic surface contains $\mathbb{Z}/5\mathbb{Z}.$ Thus, $S$ is isomorphic to the Fermat cubic surface.
Now let us prove assertion (ii). By Lemma 5.13 there are $5$ skew lines in quadratic position such that they form an orbit under the action of $\mathbb{Z}/5\mathbb{Z}\subset\mathrm{Aut}(S).$ By Corollary 5.14 we can blow them down $\mathbb{Z}/5\mathbb{Z}$-equivariantly and get the Weil restriction of scalars $Q\simeq\mathrm{R}_{\mathbf{L}/\mathbf{F}}(\mathbb{P}^{1}),$ where $\mathbf{F}\subset\mathbf{L}$ is a quadratic extension of $\mathbf{F}.$ By Remark 5.5 the blowdown gives us $5$ points in general position on $Q.$ By Proposition 3.7 we get that the set of these $5$ points is unique up to automorphism. Therefore, a smooth cubic surface such that its automorphism group contains $\mathbb{Z}/5\mathbb{Z}$ and $\Gamma\simeq\mathbb{Z}/2\mathbb{Z}$ is unique up to isomorphism. By Theorem 1.3 we get that the automorphism group of cubic surface (1.1) contains $\mathbb{Z}/5\mathbb{Z}.$ Thus, $S$ is isomorphic to cubic surface (1.1).
@@ 7. Fermat cubic surface over fields of characteristic $2$
By Theorem 1.4(i) the automorphism group of the Fermat cubic surface over $\mathbb{F}_{4^{k}}$ is isomorphic to $\mathrm{PSU}_{4}(\mathbb{F}_{2}).$ Let us briefly recall the proof of this statement, which can be found in [7, §5.1], cf. also [2, §1.3 and 2.3]. Let us consider the Hermitian bilinear form on the $4$-dimensional vector space over $\mathbb{F}_{4}$
$H\left((x,y,z,t),(x^{\prime},y^{\prime},z^{\prime},t^{\prime})\right)=x\Phi(x^{\prime})+y\Phi(y^{\prime})+z\Phi(z^{\prime})+t\Phi(t^{\prime}),$
where $\Phi:\mathbb{F}_{4}\to\mathbb{F}_{4}$ is the Frobenius automorphism defined by $\Phi(\mathtt{x})=\mathtt{x}^{2}$ for $\mathtt{x}\in\mathbb{F}_{4}.$ Therefore, since the equation of the Fermat cubic surface is
$H\left((x,y,z,t),(x,y,z,t)\right)=x^{3}+y^{3}+z^{3}+t^{3},$
we get that its automorphism group contains the group $\mathrm{PSU}_{4}(\mathbb{F}_{2}).$ So by Lemma 6.1 this means that the automorphism group of the Fermat cubic surface over $\mathbb{F}_{4}$ and algebraic extensions of $\mathbb{F}_{4}$ is isomorphic to $\mathrm{PSU}_{4}(\mathbb{F}_{2}).$
So let us consider the Fermat cubic surface over the field $\mathbb{F}_{2^{2k+1}}.$ Over its algebraic extension of degree $2$ the automorphism group of the Fermat cubic surface is isomorphic to $\mathrm{PSU}_{4}(\mathbb{F}_{2}).$ These are the elements $A$ in $\mathrm{PGL}_{4}(\mathbb{F}_{4})$ such that
(7.1) $A^{\top}\Phi(A)=E,$
where $\Phi$ acts on the entries of $A,$ $E$ is the neutral element in $\mathrm{PGL}_{4}(\mathbb{F}_{4})$ and $A^{\top}$ is the transpose of $A.$
Consider the algebraic extension $\mathbb{F}_{2^{2k+1}}\subset\mathbb{F}_{4^{2k+1}}.$ Let $\Gamma$ be the Galois group of this extension. Then the automorphism group of the smooth cubic surface over $\mathbb{F}_{2^{2k+1}}$ is a subgroup
$G=\mathrm{PSU}_{4}(\mathbb{F}_{2})^{\Gamma}\subset\mathrm{PSU}_{4}(\mathbb{F}_{2})$
of Galois invariant elements. Since the Galois group of a finite algebraic extension of finite fields is generated by the Frobenius automorphism we get that $\Gamma$ is generated by $\psi,$ i.e.
$\psi(\mathtt{x})=\mathtt{x}^{2^{2k+1}}\;\;\text{for}\;\;\mathtt{x}\in\mathbb{F}_{4^{2k+1}}.$
So we get that $G$ is a subgroup of the group $\mathrm{PGL}_{4}(\mathbb{F}_{4})$ consisting of the elements $A$ which satisfies the equation
(7.2) $A=\Phi(A).$
This means that the automorphism groups of the Fermat cubic surface over $\mathbb{F}_{2}$ and over $\mathbb{F}_{2^{2k+1}}$ are isomorphic. So it is enough to find the automorphism group of the Fermat cubic surface over $\mathbb{F}_{2}.$ By (7.1) and (7.2) these are the elements
$M\in\mathrm{GL}_{4}(\mathbb{F}_{2})$
which satisfy the equation $M^{\top}M=E.$ In other words, these are the matrices
$M=\begin{pmatrix}m_{11}&m_{12}&m_{13}&m_{14}\\ m_{21}&m_{22}&m_{23}&m_{24}\\ m_{31}&m_{32}&m_{33}&m_{34}\\ m_{41}&m_{42}&m_{43}&m_{44}\\ \end{pmatrix}\in\mathrm{GL}_{4}(\mathbb{F}_{2}),$
such that
(7.3) $\sum_{i=1}^{4}m_{ij}=1\;\;\text{for all}\;\;j=1,2,3,4\quad\text{and}$
(7.4) $\sum_{i=1}^{4}m_{ij}m_{ik}=0\;\;\text{for}\;\;j,k=1,2,3,4\;\;\text{and}\;\;j\neq k.$
Using non-degeneracy of $M,$ by (7.3) we get that any column of $M$ consists of either one $1,$ or three $1$ By (7.4) we get that if one of the column of $M$ consists of one $1,$ then the other columns also consist of one $1;$ and if one of the column of $M$ consists of three $1,$ then the other columns also consist of three $1$ The group of such elements is isomorphic to $\mathbb{Z}/2\mathbb{Z}\times\mathfrak{S}_{4}.$ This completes the proof.
@@
By Theorem 1.4(ii) the automorphism group of cubic (1.1) is isomorphic to $\mathfrak{S}_{6},$ while by Proposition 1.6 the automorphism of the Fermat cubic surface is isomorphic to $\mathbb{Z}/2\mathbb{Z}\times\mathfrak{S}_{4}.$ This proves assertion (i).
The isomorphism between (1.1) and the Fermat cubic surface over $\mathbb{F}_{4^{k}}$ is given by the collineation
@@
where $\omega\in\mathbb{F}_{4^{k}}$ is a non-trivial cube root of unity, which maps cubic surface (1.1) to the Fermat cubic surface. This proves assertion (ii).